OpenStudy (anonymous):

Can someone explain how to express -4 in trigonometric form? Really lost. Thank you!

jimthompson5910 (jim_thompson5910):

Polar form: \(\Large (r,\theta)\) Trig form: \(\Large r[\cos(\theta) + i*\sin(\theta)]\) The two forms are very similar

jimthompson5910 (jim_thompson5910):

write -4 as -4+0i that is in the form a+bi

OpenStudy (anonymous):

Ok.

jimthompson5910 (jim_thompson5910):

a = -4 b = 0 plug those values into the formulas r = sqrt(a^2 + b^2) theta = arctan(b/a)

OpenStudy (anonymous):

So r = 4 and theta = 0?

jimthompson5910 (jim_thompson5910):

r = 4 is correct but theta = 0 is not

jimthompson5910 (jim_thompson5910):



OpenStudy (anonymous):

Isn't 0/-4 equal to 0, and the arctan(0) just 0?

jimthompson5910 (jim_thompson5910):

arctan(b/a) = arctan(0/(-4)) = 0 is true but we're in the wrong spot add on 180 degrees